find the radius of the path of 35cl- ions in such a machine if v=3.0kv and b=3.0t

1
Expert's answer
2022-06-07T08:11:54-0400

"\frac{mv^2}r=qvB,\implies"

"r=\frac{mv}{qB},"

"\frac{mv^2}2=qU,\implies"

"v=\sqrt\frac{2qU}m,"

"r=\frac m{qB}\sqrt\frac{2qU}{m}=\sqrt\frac{2Um}{qB},"

"r=0.1~mm."